A piece of wire 5.5 m long is bent into a triangular shape. One side is 1.5m long and another is 2m long. Find the angles of the triangle.
Answer by amarjeeth123(569)

A piece of wire 5.5 m long is bent into a triangular shape. One side is 1.5m long and another is 2m long.The third side is 5.5-(2+1.5)=2m
The sides of the triangle are 2m,2m and 1.5m respectively.
Let the angle opposite to 1.5 m be x.
using cosine rule we get a^2=b^2+c^2-2bccosA
Substituting the values we get,
1.5^2=2^2+2^2-2(2)(2)cosx
2.25=4+4-8cosx
8cosx=8-2.25
8cosx=5.75
cosx=0.71875
x=44 degrees
Since two sides are equal two angles are equal and it is an isosceles triangle.
Let the other two angles be y
2y+44=180
2y=136
y=68
The angles of the triangle are 68,68 and 44 degrees.